OpenStudy (myrandaharris9698):

Which expression is equivalent to 15a^9b^4 over 5a^5b? Assume the denominator does not equal 0. A. 10a^4b^3 B. 3a^4b^4 C. 10a^4b^4 D. 3a^4b^3

OpenStudy (anthonyym):

\[\frac{ 15a^9b^4 }{ 5a^5b }\]

OpenStudy (anthonyym):

So do the numbers first: 15 divided by 5.

OpenStudy (myrandaharris9698):

3

OpenStudy (anthonyym):

Ok. Now \[\frac{ a^9 }{ a^5 }\] is the same as \[\frac{ a*a*a*a*a*a*a*a*a }{ a*a*a*a*a }\]. 5 of the a's cancel, and there are no more a's in the denominator. What does it become?

OpenStudy (myrandaharris9698):

3a^4

OpenStudy (anthonyym):

The property for variable to a power divided by the same variable to a power is that you subtract the powers, so in the example you subtract 9 - 5 and get a^4. Yes, now same thing with b^4 / b

OpenStudy (myrandaharris9698):

3a^4b?